The train from New Brunswick to New York Averages 60 miles per hour, while the return trip from New York to New Brunswick averages 50 miles per hour. The round trip takes 3 hours. How many minutes does it take for the train to travel from New Brunswick to New York?
Answer by mananth(16949)

distance = x
New Brunswick to New York x/ 60
New York to New Brunswick x/ 50
exrta time while uphill = 3.00 hour
x/ 50 - x/ 60 = 3.00

LCD = 7350 3000
multiply by LCD
60 - 50 x = 3.00 * 3000
10 x= 9000
/ 10
x= 900 miles
New Brunswick to New York = 900 miles